Metalic rod of 30cm of initial temperature of 42 degree celsius extended to 30.2Cm when subjected to 82 degree celsius calculate linear expansivity of the metalic material .If the material were to be used to construct an cuboid box. What is the cubic expansivity of the box.

1
Expert's answer
2022-01-27T13:24:05-0500

We know that

"V_0=0.30^3=0.027m^3"

"V=0.302^2=0.02754m^3"

We know that

"V=V_0[1+3\alpha(T-T_0)]"

Put value


"0.02754=0.027[1+3\times\alpha(82-42)]"

"1.02=1+3\times40\alpha"

"1.02-1=3\times40\alpha"

"0.02=3\times40\alpha"

"\alpha=\frac{0.02}{120}"

"\alpha=1.67\times10^{-4} /°C"